# Open-Set Deepfake Detection: A Parameter-Efficient Adaptation Method with Forgery Style Mixture

### v1

@@ IV-C Cross-dataset evaluation
Due to the significant domain gaps between training and testing datasets [59, 60] in face forgery detection, cross-dataset evaluation has been a persistent challenge. We further conduct cross-destatet evaluations to assess the generalizability of forgery detectors. We train the models on the FF++ (c23) training set (OR, DF, FF, FS, and NT) and test them on six popular Deepfake datasets, including CelebDF-v2 (CDF) [61], WildDeepfake (WDF) [62], DeepFake Detection Challenge (DFDC) [63], DeepFake Detection Challenge Preview (DFDC-P) [63], DeepForensics-1.0 (DFR) [64], and Face Forensics in the Wild (FFIW) [65]. These testing datasets simulate real-world scenarios well, encompassing various environmental variables. AUC and EER performance are reported in Table III, where ‘*’ indicates the trained model provided by the authors, and ‘†’ indicates our re-implementation using the public official code. The upper table presents the frame-level detection results, while the blue-highlighted bottom table shows video-level results. Video-level detection scores are calculated by averaging all frame scores within each video. The best and second-best results are highlighted in bold and underlined, respectively. The trainable parameter number of each detector is provided in the table for reference. Our method, OSDFD, achieves the highest average detection performance in both frame-level and video-level evaluation, with only 1.34M activated parameters. Compared to the ViT-B backbone, OSDFD consistently demonstrates significant performance improvements across all six datasets, demonstrating its exceptional generalization capability. Furthermore, OSDFD outperforms our conference version SCLoRA by a significant margin, underscoring the effectiveness of the designed CDC Adapter layer and forgery-style mixture module.
@@ IV-D Robustness evaluation
Videos circulating on social media platforms often undergo various distortions during transmission. To simulate real-world scenarios, we introduce six common perturbations: Gaussian blur, pink noise, white noise, block, brightness, and darkness. Fig. 8 showcases both the original and corresponding distorted face images. We incorporate five severity levels for each perturbation type to mimic unpredictable environments. The robustness evaluation results are depicted in Fig. 7. OSDFD does not employ data augmentations during training, facilitating efficient training, or fine-tuning for real-world applications. All models depicted in Fig. 7 are trained on the FF++ c23 training set, with distortions applied to the corresponding test set. As the severity of distortion increases, all face forgery detection methods experience performance degradation. Compared to previous arts, our proposed approach achieves superior or comparable AUC performance across all six types of perturbation. Furthermore, the average robustness results across all distortions highlight that OSDFD, with the fewest trainable parameters, demonstrates the best robustness against practical image/video perturbations.
@@
Figure 8: Input face image and its perturbated versions.
@@ IV-E1 Effectiveness of the designed modules
The proposed method showcases its adaptability by being flexibly adapted to other ViT backbones, a practical feature for researchers, engineers, and practitioners in the field of computer vision and image processing. Herein, we assess the face forgery detection performance of our OSDFD on various ViT backbones, including ViT-Tiny (ViT-T), ViT-Small (ViT-S), and ViT-Large (ViT-L). The cross-manipulation (c23) AUC results are presented in Table V. It can be readily observed that the average cross-manipulation performance improves with increasing model capacity. Nevertheless, larger ViT models tend to require significantly more computing resources and inference time, thereby impeding their real-world deployment.
In contrast, OSDFD consistently enhances the generalizability of ViT backbones across different model complexities, underscoring its practicality and adaptability. Specifically, our method achieves average AUC enhancements of 11.7%, 15.7%, and 16.7%, respectively. These results demonstrate the capability of OSDFD to seamlessly integrate into existing ViT models in a plug-and-play manner.
@@
TABLE V: Cross-Manipulation AUC performance on FF++ (c23) dataset with different ViT backbones.
Backbone OSDFD DF FF FS NT AVG
@@
IV-E2 Impacts of different ratios of mixed features
@@ IV-F Visualization results
In Fig. 11, we present the t-SNE [67] feature embedding visualization results of both the ViT-B baseline and our method OSDFD. Fig. 11 (a) and (b) showcase the results on the unseen CDF dataset, while Fig. 11 (c)-(h) depict the results on the WDF, DFDC, and FFIW datasets, respectively. Red and blue dots represent real and fake data samples in these visualizations. The manifolds of real and fake feature distributions entangle for the baseline model. This observation indicates that the baseline model struggles to discriminate forgery faces from the real ones. In contrast, OSDFD effectively separates real and fake images in the feature space. This improvement underscores the enhanced generalizability achieved by our model, attributed to the designed forgery-style mixture module and forgery-aware PEFT modules.
@@ V Conclusions and Future Work
While our proposed method is effective, we limit our scope herein to non-sequential Deepfake detection. Adapting our approach to leverage abundant sequential information could significantly improve video-level face forgery detection. The designed components can be seamlessly integrated into video transformers, achieving superior performance in uncontrolled environments. This path can be explored in future work.

### v2

@@ IV-B Cross-manipulation evaluation
Implicit ID [64] CVPR23 22.36M 83.80 24.85 - - - - 81.23 26.80 - - - - - -
SFDG [77] CVPR23 22.40M 75.83 30.30 69.27 37.70 - - 73.64 33.70 92.10 15.10 - - - -
UCF* [65] CVPR23 46.80M 77.15 30.20 77.40 29.80 69.30 35.80 68.40 37.80 80.80 27.00 68.30 36.70 73.56 32.88
LSDA [78] CVPR24 20.81M 82.70 - 71.40 - 76.30 - 71.10 - 87.50 - 84.20 - 78.87 -
DiffusionFake [66] NIPS24 19.34M 83.17 24.59 75.17 33.25 77.35 30.17 - - - - - - - -
DSRL* [79] IJCV24 128.31M 74.23 30.48 74.18 31.16 72.01 34.03 73.74 31.35 94.50 13.09 86.77 20.41 79.24 26.75
FreqNet† [80] AAAI24 1.85M 72.81 33.01 73.37 32.67 72.89 33.37 68.78 36.65 86.94 26.48 75.67 30.26 75.08 32.07
@@
UCF* [65] CVPR23 46.80M 83.70 28.20 77.20 26.10 71.00 33.50 70.90 35.80 86.40 20.40 69.70 34.20 76.48 29.70
LSDA [78] CVPR24 20.81M 89.80 - 75.60 - 81.20 - 73.50 - 89.20 - 84.20 - 82.25 -
DSRL* [79] IJCV24 128.31M 84.56 22.90 77.31 27.32 75.64 30.17 77.99 27.63 97.23 10.44 88.69 19.82 83.57 23.05
FreqNet†[80] AAAI24 1.85M 75.79 28.59 79.54 29.14 75.07 31.16 69.51 36.11 88.79 25.97 79.56 26.47 78.04 29.57
@@ IV-C Cross-dataset evaluation
Due to the significant domain gaps between training and testing datasets [83] in face forgery detection, cross-dataset evaluation has been a persistent challenge. We further conduct cross-destatet evaluations to assess the generalizability of forgery detectors. To ensure fair comparisons in cross-dataset experiments, we follow previous works [63, 76, 16] by using the FF++ (c23) dataset [22] as the training set. This dataset includes original real faces and four types of forgeries: Deepfakes, Face2Face, FaceSwap, and NeuralTextures. To balance positive and negative samples during training, real faces are augmented fourfold. Additionally, during the forgery style mixture stage, we randomly mix the four forgery types. We test the trained models on six popular Deepfake datasets, including CelebDF-v2 (CDF) [84], WildDeepfake (WDF) [25], DeepFake Detection Challenge (DFDC) [23], DeepFake Detection Challenge Preview (DFDC-P) [23], DeepForensics-1.0 (DFR) [24], and Face Forensics in the Wild (FFIW) [26]. These testing datasets simulate real-world scenarios well, encompassing various environmental variables. AUC and EER performance are reported in Table III, where ‘*’ indicates the trained model provided by the authors, and ‘†’ indicates our re-implementation using the public official code. The upper table presents the frame-level detection results, while the blue-highlighted bottom table shows video-level results. Video-level detection scores are calculated by averaging all frame scores within each video. The best and second-best results are highlighted in bold and underlined, respectively. The trainable parameter number of each detector is provided in the table for reference. Our method, OSDFD, achieves the highest average detection performance in both frame-level and video-level evaluation, with only 1.34M activated parameters. Compared to the ViT-B backbone, OSDFD consistently demonstrates significant performance improvements across all six datasets, demonstrating its exceptional generalization capability. Furthermore, OSDFD outperforms our conference version SCLoRA by a significant margin, underscoring the effectiveness of the designed CDC Adapter layer and forgery-style mixture module.
@@ IV-D Robustness evaluation
We incorporate five severity levels for each perturbation type to mimic unpredictable environments. The robustness evaluation results are depicted in Fig. 8. OSDFD does not employ data augmentations during training, facilitating efficient training, or fine-tuning for real-world applications. All models depicted in Fig. 8 are trained on the FF++ c23 training set, with distortions applied to the corresponding test set. As the severity of distortion increases, all face forgery detection methods experience performance degradation. Compared to previous arts, our proposed approach achieves superior or comparable AUC performance across all six types of perturbation. Furthermore, the average robustness results across all distortions highlight that OSDFD, with the fewest trainable parameters, demonstrates the best robustness against practical image/video perturbations. Our proposed style mixture module contributes to enhancing model robustness by increasing the diversity of forgery domains, thereby reducing the distribution gap between clean training data and corrupted testing data.
@@ IV-E1 Effectiveness of the designed modules
The proposed method showcases its adaptability by being flexibly adapted to other transformer backbones. OSDFD comprises forgery-aware, parameter-efficient fine-tuning modules and a forgery-style mixture module, both designed as plug-and-play components that improve the generalization capability of transformer backbones. As shown in Table V, we apply our method OSDFD to various models, including ViT [17] and Swin [85]. The results demonstrate that OSDFD consistently enhances performance across all backbones, leading to significant improvements in AUC scores in the cross-manipulation setting. Notably, the highest performance gains are observed in ViT-L (+0.167 AVG AUC) and Swin-L (+0.112 AVG AUC), suggesting that OSDFD is particularly effective when integrated with larger backbones. While the average cross-manipulation performance improves with increasing model capacity, larger models tend to require significantly more computing resources and inference time, thereby impeding their real-world deployment. These findings indicate that OSDFD is a flexible and effective module for improving the generalizability of transformer-based face forgery detection models.
@@
TABLE V: Cross-manipulation AUC performance on FF++ (c23) dataset with different backbones.
Backbone OSDFD DF $F$ FS NT AVG
@@
Swin-S - 0.897 0.768 0.400 0.666 0.683
Swin-S 0.922 0.867 0.538 0.759 0.772
Swin-B - 0.877 0.715 0.463 0.668 0.681
Swin-B 0.943 0.888 0.583 0.734 0.787
Swin-L - 0.923 0.743 0.446 0.646 0.690
Swin-L 0.956 0.906 0.601 0.745 0.802
@@
Figure 10: Cross-dataset evaluation results using different seeds: (a) AUC; (b) EER.
@@
IV-E2 Effectiveness of FSM
To further evaluate the effectiveness of the FSM module, we report additional metrics, including mAP, ACC, and FNR, in Table VII. The cross-dataset evaluation results with and without FSM are presented in Table VII, showing that the inclusion of FSM consistently improves face forgery detection performance. Notably, FSM significantly enhances the FNR, indicating that it effectively mitigates forgery domain gaps between training and testing sets. This, in turn, enhances the model’s ability to detect forgery faces from another point of view.
IV-E3 Effectiveness of CDC
To validate the effectiveness of the CDC operator, we conduct an ablation study by integrating various conventional convolutions with different kernel sizes. Table VIII presents the model’s cross-dataset experimental results using different local feature extractors, where Conv3×3, Conv5×5, and Conv7×7 represent standard convolutions with kernel sizes of $3\times 3$, $5\times 5$, and $7\times 7$, respectively. The Second-order Difference Convolution (SDC) module, as proposed by Fei et al. [54], employs a second-order local anomaly learning mechanism to detect anomalies in local regions. These baseline results are obtained by replacing the CDC convolution in Fig. 4 with the corresponding operators. The results demonstrate that the CDC operator consistently outperforms other local feature extractors across various unforeseen datasets, which can be attributed to its outstanding capability for local high-frequency feature extraction.
IV-E4 Impacts of different ratios of mixed features
@@ IV-F1 T-SNE feature map visualization
In Fig. 12, we present the t-SNE [87] feature embedding visualization results of both the ViT-B baseline and our method OSDFD. Fig. 12 (a) and (b) showcase the results on the unseen CDF dataset, while Fig. 12 (c)-(h) depict the results on the WDF, DFDC, and FFIW datasets, respectively. Red and blue dots represent real and fake data samples in these visualizations. The manifolds of real and fake feature distributions entangle for the baseline model. This observation indicates that the baseline model struggles to discriminate forgery faces from the real ones. In contrast, OSDFD effectively separates real and fake images in the feature space. Specifically, for CDF and FFIW datasets, it is obvious that OSDFD can better discriminate real faces from fake ones compared to the ViT baseline. The AUC score goes from 72.35 to 83.35 on the CDF dataset, and 68.49 to 84.61 on the FFIW dataset. In turn, ViT-B can achieve good detection performance on WDF and DFDC datasets, while our proposed method OSDFD, benefiting from the designed forgery-aware parameter efficient fine-tuning and the forgery style mixture modules, exhibits a better face forgery detection performance.
@@
TABLE VII: Impacts of Forgery Style Mixture (FSM).
Setting CDF DFR WDF
[…]
w/o FSM 88.91 72.87 30.69 93.42 84.28 15.40 84.74 70.94 23.61
w/ FSM 90.33 75.05 22.82 95.25 88.22 11.01 86.70 72.77 16.13
@@ IV-F2 Grad-CAM map visualization
TABLE VIII: Comparison with other local-frequency extractor.
Setting CDF DFR WDF AVG
[…]
Conv3×3 82.26 26.66 94.57 11.59 78.00 29.20 84.94 22.48
Conv5×5 80.39 27.32 91.35 15.75 76.98 30.18 82.91 24.42
Conv7×7 80.44 27.71 92.92 14.25 78.10 29.09 83.82 23.68
SDC [54] 79.14 27.89 92.87 11.21 77.72 28.32 83.24 22.47
CDC 83.35 24.47 95.54 9.90 78.53 28.46 85.81 20.94
IV-F3 Cross-dataset results using different seeds
To mitigate the influence of different seeds, we train our models for seven trials using seeds 0-6. The cross-dataset AUC and EER performances are shown in Fig. 10 (a) and (b), respectively. The results indicate that our approach consistently achieves promising performance regardless of the different seeds, demonstrating its robustness across varied initializations. This robustness further underscores the superiority of our method from an additional perspective.
@@ V Conclusions and Future Work
While our proposed method is effective, we limit our scope herein to non-sequential Deepfake detection. Adapting our approach to leverage abundant sequential information could significantly improve video-level face forgery detection. The designed components can be seamlessly integrated into video transformers. Extending our designed method to video transformers for video-level Deepfake detection opens an important research path forward. We anticipate that the proposed modules can well adapt to video ViTs and enhance the generalizability. Furthermore, our method exhibits strong robustness against common image perturb ations, as shown in Fig. 8. However, its performance degrades under the highest perturbation severity (level 5). Additionally, although our approach contributes to safeguarding digital media integrity, Deepfake detection models may exhibit bias across different demographic groups due to imbalances in the training dataset. In future work, our aim is to improve robustness and ensure fairness in Deepfake detection.